# Spectral properties of two-dimensional half-space semi-Dirac semi-metals

### v1

@@ 4 The spectral properties in semi-Dirac semi-metals
$4(\lambda_{1}+\delta)^{2}-4\lambda_{1}-8\lambda_{1}\delta-4\lambda_{1}^{2}=4\delta^{2}-4\lambda_{1}.$
If $\lambda_{1}=\frac{\pi^{2}}{(b-a)^{2}}<\delta^{2}$ then the corresponding quadratic equation has two solutions $V_{1},V_{2}:V_{1}<V_{2}$. Therefore, for any $V\in(V_{1},V_{2})$ in $[a,b]^{2}$ we have $\|Hv\|^{2}-\delta^{2}\|v\|^{2}<0$. It deduces the existence of the discrete spectrum of $H^{2}$ as well as $H$. It concludes the proof.
@@
$g_{n}(x,y)=\left\{\begin{array}[]{rcl}&1&\mbox{ if }\sqrt{x^{2}+y^{2}}\leq n\,,\\ &g(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})&\mbox{ if }n<\sqrt{x^{2}+y^{2}}<n^{2}\,,\\ &0&\mbox{ if }\sqrt{x^{2}+y^{2}}\geq n^{2}\,,\\ \end{array}\right.$ (30)
@@
we consider the existence of the discrete spectrum not lying in the essential spectrum of $H_{\epsilon}$ depending on the sign of $A_{\epsilon}$.
@@
Let $w_{ij}\in L^{1}(\mathbb{R}^{2}_{+})\cap L^{\infty}(\mathbb{R}^{2}_{+})$ for each $i,j=1,2$ and vanishing at infinity. If $A_{\epsilon}<0$ then the discrete spectrum in $-(\delta,\delta)$ of $H_{\epsilon}$ is non-empty.
@@
$\quad+\|(-\partial^{2}_{x}+\delta+\epsilon w_{21})g_{n}+(i\partial_{y}+\epsilon w_{22})g_{n}\|^{2}\,,$
@@
If $A_{\epsilon}<0$ then there exist $n*\in\mathbb{N},B_{\epsilon}<0$ such that for all $*n\geq n*$ : $\|H\psi_{n}\|^{2}-\delta^{2}\|\psi_{n}\|^{2}<B_{\epsilon}<0$. As a result, the discrete spectrum of $H_{\epsilon}$ is non-empty or there exists at least one eigenvalue in $(-\delta,\delta)$ of $H_{\epsilon}$. It concludes the proof.
@@
Another comment referring to the case $w_{11}=w_{22}=0,w_{12}=w_{21}\geq 0$ is if there exists the discrete spectrum of $H_{\epsilon}$ then that has to be embedded eigenvalues lying in the essential spectrum of $H_{\epsilon}$. This result is derived by using the analogous arguments given in [7].
@@
It is noteworthy that the sufficient conditions posed on $\Re w_{12}$ in Corollary 1 and in Theorem 6 are not identical. The former relies on the integral of $\Re w_{12}$ in +$\mathbb{R}^{2}_{+}$, while the latter requires rigorous values of $V=\Re w_{12}$ in a certain range $[a,b]^{2}\in\mathbb{R}^{2}_{+}$.
@@ Acknowledgment
The author was partially supported by the EXPRO grant No. 20-17749X of the Czech Science Foundation.
Declarations
Conflict of interest There is no conflict of interest to declare.

### v2

@@ 4 The spectral properties of semi-Dirac semi-metals
$4(-\frac{9}{5}\lambda_{1}+\delta)^{2}-4\lambda_{1}+\frac{72}{5}\lambda_{1}\delta-36\lambda_{1}^{2}=4\delta^{2}-4\lambda_{1}-\frac{576}{25}\lambda_{1}^{2}.$
If
$\lambda_{1}+\frac{576}{100}\,\lambda_{1}^{2}<\frac{\pi^{2}}{(b-a)^{2}}+6\,\frac{\pi^{4}}{(b-a)^{4}}<\delta^{2}\,$ (35)
then the corresponding quadratic equation has two solutions $V_{1},V_{2}:V_{1}<V_{2}$. Therefore, for any $V\in(V_{1},V_{2})$ in $[a,b]^{2}$ we have $\|Hv\|^{2}-\delta^{2}\|v\|^{2}<0$. It deduces the existence of the discrete spectrum of $H^{2}$ as well as $H$. It concludes the proof.
This result provides us a local selection for the potential $V$ defined in an arbitrary square in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ with the aim of establishing the existence of the discrete spectrum of the semi-Dirac semi-Laplacian problem in the half-space. It can be seen that the rigorous relation between $\delta$ and $b-a$ described in (35) can be supplemented by a numerical evidence. Figure 2 illustrates the validity region of (35).
Figure 2: The blue colour indicates the validity region of (35).
Remark 2.
Based on our analysis, it is apparent that the local establishment for $V$ can be alternatively determined in an appropriate rectangle in +$\mathbb{R}^{2}_{+}$ by virtue of employing the separation of variables for trial functions. Nevertheless, the setting of that in squares in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ perhaps completely fulfills our principle purposes.
@@
$g_{n}(x,y)=\left\{\begin{array}[]{rcl}&1&\mbox{ if }\sqrt{x^{2}+y^{2}}\leq n\,,\\ &g(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})&\mbox{ if }n<\sqrt{x^{2}+y^{2}}<n^{2}\,,\\ &0&\mbox{ if }\sqrt{x^{2}+y^{2}}\geq n^{2}\,.\\ \end{array}\right.$ (36)
@@
we consider the existence of the discrete spectrum of $H_{\epsilon}$ depending on the sign of $A_{\epsilon}$.
@@
Let $w_{ij}\in L^{1}(\mathbb{R}^{2}_{+})\cap L^{\infty}(\mathbb{R}^{2}_{+}),w_{ij}=\overline{w_{ji}}$ for $i,j=1,2$ and vanishing at infinity. If $A_{\epsilon}<0$ then the discrete spectrum of $H_{\epsilon}$ is non-empty.
@@
$\quad+\|(-\partial^{2}_{x}+\delta+\epsilon w_{21})g_{n}+(i\partial_{y}+\epsilon w_{22})g_{n}\|^{2}-2\delta^{2}\|g_{n}\|^{2}\,,$
@@
If $A_{\epsilon}<0$ then there exist $n*\in\mathbb{N},B_{\epsilon}<0$ such that for all $n\geq n*$ : $\|H\psi_{n}\|^{2}-\delta^{2}\|\psi_{n}\|^{2}<B_{\epsilon}<0$. As a result, the discrete spectrum of $H_{\epsilon}$ is non-empty or there exists at least one eigenvalue in $(-\delta,\delta)$ of $H_{\epsilon}$. It concludes the proof.
@@
Another comment referring to the case $w_{11}=w_{22}=0,\mathbb{R}\ni w_{12}=w_{21}\geq 0$ is if there exists the discrete spectrum of $H_{\epsilon}$ then that has to be embedded eigenvalues lying in the essential spectrum of $H_{\epsilon}$. Consequently, the discrete spectrum of the operator is empty in this case. This result is derived by using the analogous arguments given in [7].
@@
It is noteworthy that the sufficient conditions posed on $\Re w_{12}$ in Corollary 1 and in Theorem 6 are not identical. The former relies on the integral of $\Re w_{12}$ in +$\mathbb{R}^{2}_{+}$, while the latter requires rigorous values of $V=\Re w_{12}$ in a certain range $[a,b]^{2}\subset\subset\mathbb{R}^{2}_{+}$.
Remark 3.
The spectral properties of the semi-Dirac semi-Laplacian are unstable and mismatched if we alter the half space +$\mathbb{R}^{2}_{+}$ into the other one $(0,\infty)\times\mathbb{R}$ due to the fact that the corresponding semi-Dirac operator with Dirichlet boundary conditions is also closed, symmetric but not self-adjoint in the half space $(0,\infty)\times\mathbb{R}$. It can be seen as a distinction between that of the semi-Dirac semi-metals and Dirac operators under some specific sorts of boundary conditions.
@@ Acknowledgment
The author is grateful to D. Krejčiřík for useful comments and was partially supported by the EXPRO grant No. 20-17749X of the Czech Science Foundation.